OpenStudy (anonymous):

If the slower ship in the preceding exercise leaves at noon and the other at 1p.m. , how far apart are they at 2p.m....................... anyone can help me i will give medals and i will be your fan

OpenStudy (anonymous):

@uri

OpenStudy (anonymous):

@CaseyCarns

OpenStudy (anonymous):

@mathmath333

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

buddy can you help me here

OpenStudy (mathmath333):

is the question complete

OpenStudy (anonymous):

yep

OpenStudy (mathmath333):

i think the question is lacking some information

OpenStudy (anonymous):

wait i will check again

OpenStudy (anonymous):

can i give another question here?

OpenStudy (mathmath333):

yes u caan

OpenStudy (anonymous):

Point A and B are separated by an obstacle. In order to find the distance between them, a third point C is selected which is 120 yards from A and 150 yards from B. The angle ACB is measured to 80degrees 10 min. Find the distance from A to B. CAN YOU USE THE COSINE LAW...AND CAN U DRAW THE FIGURE BUDDY :D

OpenStudy (mathmath333):



OpenStudy (anonymous):



OpenStudy (anonymous):

mr. math math can you show me the solution?

OpenStudy (mathmath333):

law of cosines is used here http://www.mathsisfun.com/algebra/trig-cosine-law.html

OpenStudy (anonymous):

tnx

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{cosC=\dfrac{AC^2+BC^2-AB^2}{2\times AB\times BC }\\~\\ cos(80+\frac{1}{6})=\dfrac{120^2+150^2-AB^2}{2\times 150\times 120}\\~\\ cos(80+\frac{1}{6})\times 2\times 150\times 120=120^2+150^2-AB^2\\~\\ AB^2=120^2+150^2- cos(80+\frac{1}{6})\times 2\times 150\times 120\\~\\ AB \approx \pm 175}\end{align}\)